# ConfuGuard: A Low False-Positive System for Detecting Malicious and Stealthy Typosquatting Threats in Package Registries

### v2

@@ 6.6 EQ5: System Accuracy and Efficiency
Method
We evaluated the end-to-end system accuracy and efficiency (i.e., latency and throughput) by running ConfuGuard and Typomind on NeupaneDB and ConfuDB.
Result
Table 6 presents the accuracy measurements. Across the datasets, ConfuGuard generally exhibits stronger performance in detecting active threats, achieving higher recall than Typomind while both maintain perfect precision. In the case of stealthy threats, the results are mixed: Typomind excels in the NeupaneDB dataset with a perfect recall, yet its performance declines in ConfuDB where ConfuGuard outperforms it. For benign cases, the accuracy differs notably, as Typomind misclassifies benign packages in NeupaneDB, whereas it performs better on ConfuDB, though still lower than ConfuGuard’s performance in that category. Overall, ConfuGuard demonstrates strengths in detecting active and stealthy threats across datasets, reducing the FPR from 77% (section 4.2) to 13% (72/543).
Table 7: System latency and throughput measurements. ConfuGuard has higher latency due to the additional benignity filter. Our industry partner considers this performance overhead acceptable.
ConfuGuard Typomind
Latency (ms/pkg) 6816 120
Throughput (#pkgs/s) 0.08 0.26
Table 7 indicates that while ConfuGuard incurs significantly higher latency (6816 ms per package vs. 120 ms for Typomind) and lower throughput (0.08 vs. 0.26 packages/s), these trade-offs are acceptable in production because the additional latency—stemming from benignity filtering and reliance on LLM-based metadata verification—substantially enhances detection accuracy by reducing false positives. Despite slower processing due to dependencies on the OpenAI API and o3-mini, the system scales effectively across large registries, and future work may optimize inference times without compromising accuracy. Notably, ConfuGuard significantly reduces the workload on analysts, resulting in notable cost savings.
@@ 7 Discussion
7.1 Three Lessons Learned from Production
Lesson 1: ConfuGuard prevents real confusion attacks
Over the past two months, we deployed the system in our industry partner’s production environment, during which we identified and confirmed 301 threats, with 37,266 additional threats under review.
Lesson 2: False Positives Harm Our Customers.
Table 3 represents the most recent metadata verification rules. In deployment, we found additional cases where the package has a very similar name and the README of their package was missing which made our system classify it as suspicious stealth attack. That package was owned by one of our partner’s customers and served as a “transitive package”. The customer felt that being flagged harmed their reputation. We added an allow-list tied to our partner’s customers.
Lesson 3: Ontology Matters.
ConfuGuard is deployed as part of a security analysis pipeline, with human analysts to triage the results (section 5.6). Analysts need to be able to interpret the results of ConfuGuard. The analysts feel that the package confusion categories outlined in Table 2 do not provide sufficient explanatory power. We propose to refine the confusion taxonomy to consider the malicious content and intent behind each package, including a risk-level classification.
7.2 Limitations and Security Analysis
This section discusses our system limitations and how attackers might bypass ConfuGuard.
Gaming Metrics
@@ 7.2 Limitations and Security Analysis
Our system relies on software metrics to gauge the likelihood that a package is a confusion attack. These metrics might be gamed. There has been little formal study of the feasibility of gaming these metrics, but recent work suggests both the possibility and some real-world examples [24].
Limitations in Neighbor Search
@@
One significant limitation of ConfuGuard is its inability to handle short names or acronyms. FastText struggles with short words (e.g., xml vs. yml have a similarity score of 0.7). The model’s reliance on character n-grams often fails to capture subtle similarities effectively in such cases, providing an avenue for attackers to exploit short package names. To mitigate this, we implemented a list of potential substitutions to identify cases of visual or phonetic ambiguity, and we combine embedding similarity with Levenshtein distance. This hybrid approach improves neighbor search for short names, but increases the computational cost does not fully resolve the concern.
Bypassing Metadata Verification
@@
Using an LLM for benign filtering introduces correctness risks (hallucination, jailbreaking [66]). Adversaries might exploit this by tailoring their package name or metadata to persuade the LLM that the package is trustworthy using techniques like prompt injection or model hijacking [32, 69].
@@ 7.3 Future Directions
Enhancing Representations for Package Confusion Detection:
Improving the representation of package names is crucial for more robust detection. While FastText captures semantic similarity through subword embeddings, it struggles with typographical variations, particularly for short words or acronyms. Fine-tuning FastText or training a more efficient model on domain-specific corpora including both correct and misspelled terms can address these limitations. Augmenting training data with synthetic typos and incorporating typo normalization or correction techniques before embedding generation can significantly reduce errors. Advanced models, such as transformer-based architectures fine-tuned with contrastive learning on typo-specific datasets, present a promising alternative for enhancing detection accuracy and reliability. This approach has proven effective in combating domain typosquatting, but no research has been conducted targeting package typosquatting [koide2023phishreplicant]. One challenge is the limited availability of data for verifying package squatting cases. LLMs might help here [59].
Mitigating LLM Hallucination in Code Generation
The increasing use of LLMs for code generation has introduced new challenges, as these models often hallucinate package names or generate commands for nonexistent or maliciously similar packages [51]. These hallucinations pose serious threats to the security of the software supply chain [64, 59]. Addressing this issue requires implementing typo or hallucination correction mechanisms in LLM-based package recommenders. Verifying package legitimacy, detecting typos, and integrating contextual checks can prevent the propagation of incorrect package names, reducing the risks associated with hallucinations.
Meta-Learning for Malicious Package Detection
Meta-learning approaches offer significant potential for improving malicious package detection. By leveraging anomaly detection techniques and metadata analysis [23], systems can dynamically adapt to evolving attack strategies. Meta-learning frameworks could analyze patterns across registries and rapidly identify emerging threats, enhancing the scalability and robustness of detection systems. Integrating such frameworks will be key to staying ahead of increasingly sophisticated attackers.
@@ 8 Conclusion
We present ConfuGuard, an embedding-based package confusion detection system. Based on real-world attack patterns, we refined the package confusion definition and developed a taxonomy based on engineering practices. ConfuGuard is being used in production at our industry partner and contributed to 301 confirmed package confusion threats in three months. Compared to SOTA methods, our system is good at capturing additional confusion categories, achieves a substantially lower false-positive rate, and maintains acceptable latency, making it well-suited for deployment on SPR backends while remaining effective for frontend on-demand requests. We shared our insights from production experience, customer feedback, the need for an improved ontology, and outlined future directions.
Data Availability and Research Ethics
Our artifact is available at: https://github.com/confuguard/confuguard. For replicability, we include all results and the code of the system, except the commercial metadata database (section 5.1).
Our work poses limited ethical concerns. We analyzed public packages and disclosed suspicious ones to our industry partner.

### v3

@@ 6.6 EQ5: System Accuracy and Efficiency
[…]
Result Table˜5 presents the accuracy measurements. Across the datasets, ConfuGuard generally exhibits stronger performance in detecting active threats, achieving higher recall than Typomind while both maintain perfect precision. In the case of stealthy threats, the results are mixed: Typomind excels in the NeupaneDB dataset with a perfect recall, yet its relative performance declines on ConfuDB. For benign cases, the accuracy differs notably, as Typomind misclassifies benign packages in NeupaneDB, whereas it performs better on ConfuDB, though still lower than ConfuGuard’s performance in that category. Overall, ConfuGuard demonstrates strengths in detecting active and stealthy threats across datasets, reducing the FPR from 80% (section˜4.2) to 28% (133/480).
[…]
@@ 7 Discussion
7.1 Improvements on Prior Detectors
Having described ConfuGuard’s design in section˜5 and quantified its performance in section˜6, we now summarize ConfuGuard’s four key advances over previous detectors such as Typomind and SpellBound:
  1. 1.
    Hierarchy-aware detection (section˜5.3). By splitting and scoring each namespace component, ConfuGuard can detect compound and impersonation squats in hierarchical naming ecosystems. Previous detectors only treated flat naming ecosystems, reducing their ability to identify the attack strategies of impersonation squatting, compound squatting, and domain confusion (section˜6.5).
  2. 2.
    Syntactic-Semantic Matching at Scale (section˜5.4). ConfuGuard leverages a fine-tuned FastText model to identify potential package confusions with high recall and low latency, avoiding the exhaustive comparisons typical of prior tools (section˜6.2, section˜6.3).
  3. 3.
    Metadata-driven benignity filter (section˜5.5). ConfuGuard incorporates 15 heuristics based on package metadata to accurately identify package confusions. This technique reduces the false positive rate by 52%, compared to prior approaches (section˜6.4, section˜6.5).
  4. 4.
    Engineered for production deployment (section˜5.6). ConfuGuard can be integrated into existing security workflows, such as those used by our industry partner. It detects package confusion threats at the scale required by modern software package repositories — which prior tools cannot achieve (section˜6.6). In practice, ConfuGuard reduces manual review effort and cost for security analysts by over 50%, so that even with higher per-package processing time, operational efficiency improves substantially.
7.2 Limitations
@@
The primary design limitation of ConfuGuard is its reliance on software metrics to gauge the likelihood that a package is a confusion attack. These metrics might be gamed. There has been little formal study of this risk, but recent work suggests both the possibility and some real-world examples [30, 34].
@@
A second design limitation is that the neighbor search relies on sub-word embeddings, which struggle to distinguish acronyms and short names (curse of dimensionality). FastText struggles with short words (e.g., xml vs. yml have a similarity score of 0.7). To mitigate attacks on short package names, we implemented a list of potential substitutions to identify cases of visual or phonetic ambiguity, and we combine embedding similarity with Levenshtein distance.
@@
In our implementation, we rely on an LLM for benign filtering. This introduces correctness risks (hallucination, jailbreaking [78]). Adversaries might exploit this by tailoring their package name or metadata to persuade the LLM that the package is trustworthy using techniques like prompt injection or model hijacking [41, 81].
7.3 Lessons Learned from Production
We share four lessons learned over the six months of production deployment of ConfuGuard.
Lesson 1: ConfuGuard prevents real confusion attacks. Over the past six months, we deployed the system in our industry partner’s production environment, during which we identified and confirmed 630 threats, with 50,069 additional threats under review. Based on the false positive rate in our analysis so far, we estimate that among these threats, there are $\sim$17,500 real attacks.
Lesson 2: Simplified LLM Prompts Unlock More Robust Reasoning.
Our production experience revealed that overly detailed, rule-by-rule prompts can constrain an LLM’s ability to synthesize diverse signals, leading to both blind spots and bias [27]. For example, by instructing the model to treat any package with a namespace as “likely fork,” we missed malicious variants that reused legitimate namespaces but embedded suspicious payloads in their README or metadata. We replaced the myriad micro-rules with a decision framework approach (section˜5.5.2), so the LLM can weigh name similarity, maintainer reputation, and other signals together.
Lesson 3: Ontology Matters. ConfuGuard is deployed as part of a security analysis pipeline, with human analysts to triage the results (section˜5.6). Analysts must interpret the results of ConfuGuard, and feel the package confusion categories outlined in Table˜2 do not provide sufficient explanatory power. We propose to refine the confusion taxonomy to consider the malicious content and intent behind each package, including a risk-level classification.
Lesson 4: Large-Scale Empirical Feedback is Crucial. Our work with ConfuGuard and the experience-driven optimization pipeline (section˜5.6) underscores that only through large-scale empirical measurement can we systematically uncover the bottlenecks, failure modes, and hard edge cases that theory alone cannot predict. By continuously injecting human-in-the-loop analysis — where experts inspect anomalies and feed targeted insights back into the design — we close the gap between observed behavior and ideal performance by turning raw operational data into actionable system improvements.
Lesson 5: False Positives Harm Our Customers. Table˜3 represents the most recent benignity check rules. In deployment, we found additional cases where a package had a very similar name and no README, which made our system classify it as a suspicious stealth attack. That package was owned by one of our partner’s customers and served as a “transitive package”. The customer felt that being flagged harmed their reputation. To address such cases, where engineering convenience conflicts with security [26] (cf. section˜2.4.2), we added an allow-list of our partner’s customers.
@@ 7.4 Future Directions
Enhancing Representations. Improving the representation of package names is crucial for more robust detection. While FastText captures semantic similarity through subword embeddings, it struggles with typographical variations, particularly for short words or acronyms. In our experiments on Hugging Face packages, we observed uniformly high similarity scores due to long, acronym-dense names and extensive forking activity, which undermines standard confusion-detection methods [31]. LLMs can generate rich, domain-specific corpora, complete with synthetic typos and acronym variations. Applying (semi-)supervised learning to such tailored datasets can address these limitations.
One challenge based on our production experience is the limited availability of data for verifying package confusions. LLMs might help by augmenting training data with synthetic typos [71]. Advanced models, such as transformer-based architectures fine-tuned with contrastive learning on such typo-specific datasets, present a promising way for better detection accuracy and reliability. This approach has proven effective in combating domain typosquatting, but no research has been conducted targeting package confusions [36].
Human-in-the-Loop Optimizations. Our deployed system offers opportunities for further optimization, particularly through the incorporation of additional human-annotated ground truth. Drawing inspiration from human-in-the-loop reinforcement learning — where domain experts iteratively guide and refine agent behaviors — and self-adaptive prompting [76, 57], we propose integrating real-time feedback from security analysts into the reasoning processes of LLMs [25]. Analysts’ corrections and augmentations of intermediate model outputs serve as dynamic expert feedback, enabling improvements in both the prompts and the model’s behavior.
Anomaly Detection for Malicious Package Detection. Anomaly detection approaches offer significant potential for improving malicious package detection. By leveraging anomaly detection techniques and metadata analysis [29], future package confusion detectors might dynamically adapt to evolving attack strategies, e.g., learning new benignity rules without human oversight.
@@ 8 Conclusion
Modern software applications integrate third-party packages. Attackers therefore publish packages with names deliberately intended to confuse engineers, seeking to inject malware into production systems. Accurately detecting these attacks at ecosystem scale requires a low-latency system that accounts for human interpretations of package names. We showed that these requirements can be met by an embedding-based detection system combined with heuristics based on real-world attack patterns. Compared to the state-of-the-art, our ConfuGuard system captures additional confusion categories, achieves a substantially lower false-positive rate, and maintains acceptable latency for use either on registry backends or in handling on-demand requests. We conclude with insights from production experience and customer feedback.
Data Availability
Our artifact is available at: https://github.com/confuguard/confuguard. It includes everything but the commercial metadata database.